# Implementation Study of Cost-Effective Verification for Pietrzak’s Verifiable Delay Function in Ethereum Smart Contracts

### v3

@@ 3.3 Environment for Gas Usage Measurement
Here we describe the setup of this research used to measure gas usage in Ethereum smart contracts, detailing the use of Solidity and the related framework.
3.3.1 Development Environment and Compiler Version
The development of the contract was conducted using Solidity version 0.8.23. This choice was driven by the compatibility and full support provided by the Hardhat framework, a comprehensive Ethereum development environment with integrated debugging capabilities. At the start of the development process, Solidity 0.8.23 was the latest compiler version fully supported by Hardhat for debugging and other essential development functionalities. Generally, gas optimization is best achieved with the latest version of the compiler.
@@ 3.3.2 Optimizer Configuration
For configuration, the core parameter of the Solidity optimizer is Runs which indicates how often each opcode will be used in the deployed contract. That is, higher Runs implies less gas cost for execution and higher gas cost for deployment. We set the optimizer parameter Runs to 4,294,967,295($2^{32}-1$), which is the maximum value of the parameter, indicating the highest level of optimization for execution.
@@
In this step, we can enhance the Yul optimizer’s efficiency. We replaced the msize() operation with the free memory pointer. This adjustment necessitated manual updates to the memory indexes, a process rigorously tested to ensure the integrity of our implementation. We also applied memory-safe assembly blocks in our code. This change not only aligns with Solidity’s memory model but also activates Via-IR in the compiler settings, resulting in significant gas savings.
3.3.3 Gas Measurement
We implemented external functions in MinimalApplication.sol, the major source file for implementation, to measure the gas used with different strategies. The gas cost depends on the order of the function selectors in these external functions, but the differences are considered to be negligible, with a maximum difference of about 60 gas used across the seven external functions. Therefore, an error margin of about 60 gas exists for the results presented in the figures.
3.3.4 Local Testing Environment
We utilized the Hardhat node, provided by the Hardhat framework, to create a local testing environment. The amount of gas usage was obtained using the eth_estimateGas method, which generates and returns an estimate of the gas required to complete the transaction. This method accurately measures the gas used in the execution of different contract functions and provides valuable insights into the efficiency of the implementations. It is important to note that the actual gas cost incurred on the real network is determined by multiplying the estimated gasUsed by the gasPrice at the Ethereum network circumstance.
4 Data and Accompanying Gas Cost
@@ 4.1 Data Structure
In this study, the VDFs are implemented with key sizes of 2048 and 3072 bits. However, the EVM natively supports arithmetic operators only up to 256-bit numbers. Therefore, to facilitate efficient computation of numbers exceeding 256 bits, it is necessary to employ customized data formats. We revised the big number library implementation for Solidity developed by Firo [9]. The original data structure in this library includes bytes data (bytes val), bit length (uint bitlen), and a sign indicator (bool neg). Among the three elements, we deleted the sign indicator and its related low-level logic since our implementation does not require handling negative values. This library optimizes on-chain operations such as addition and subtraction by adding leading zeros for 32-byte alignment with the EVM’s memory word size, eliminating offset management. Also, it utilizes the bit length data field to optimize multiplication and comparison operations. Therefore, we format the bytes data and pre-calculate bit length off-chain before submission, significantly reducing unnecessary on-chain computations. For example, the data format for a 2048-bit representation in JSON is as shown in Listing 1. In this way, we can reduce the gas cost even though the calldata size increases. This decision is reasonable because the primary goal is to reduce the gas cost, not the transaction calldata size.
@@
4.2 Gas Cost of Calldata
The gas cost for Ethereum transaction calldata was fixed in EIP-2028 "Transaction data gas cost reduction" [10]. According to the document, the cost of calldata is divided into two cases:
  • Non-zero byte: 16 gas
  • Zero byte: 4 gas
@@ 4.2 Gas Cost of Calldata
In our implementation, the function verifyRecursiveHalvingProof processes the generalized verification (Algorithm 4). It is defined as the Listing 2 in Solidity codes.
1function verifyRecursiveHalvingProof(
2 BigNumber[] memory v,
3 BigNumber memory x,
4 BigNumber memory y,
5 BigNumber memory n,
6 uint256 delta,
7 uint256 T
8)
Listing 2 Interface of the Halving Verification Function
A call to this function involves encoding parameter values. Start by computing the keccak256 of the UTF-8 byte representation of the string111verifyRecursiveHalvingProof((bytes,uint256)[],(bytes,uint256),(bytes,uint256),(bytes,uint256),uint256,uint256) based on Listing 2 and use the first four bytes, which results in 0xd8e6ac60.
@@ 5 BigNumber memory n,
Considering the precision of data and its cost, we performed experiments to test the size of calldata and the related gas usage. The related gas usage comprises of the intrinsic gas and data dispatching gas. The intrinsic gas in EVM refers the minimum cost for a transaction. It contains the constant transaction cost (21000 gas) and the cost for data supply. For verification, we need not only to allocate data in a transaction, but also to dispatch data into the verification function. The dispatching cost is proportional to the data size. Therefore, we decided to include the dispatching cost into the data cost in this study. Figure 2(a) for $\lambda=2048$ and 2(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage. We can see the gas usage and the calldata size simply and linearly increase with the number of proofs.
Figure 2: Experiment results of the calldata length and cost for the proofs
(a) Calldata size and gas cost including intrinsic gas and dispatching gas ($\lambda=2048$)
(b) Calldata size and gas cost including intrinsic gas and dispatching gas ($\lambda=3072$)
Finally, to elucidate the relationship between transaction size and gas costs, we perform a regression analysis. By applying the least squares method to approximate the best fit, we derive the following linear regression equations 3 and 4 for data sizes of 2048 and 3072 bytes, respectively:
@@
5 Arithmetic Operations and Accompanying Gas Cost
This section focuses on analyzing the expenses associated with performing arithmetic operations in the context of verifying the Pietrzak VDF. The cost consists of two components: the cost of repeated halving and the cost of modular exponentiation for the final check. The algorithm executes a halving operation and then proceeds with a modular exponentiation verification. The modular exponentiation cost is studied before the halving cost, as the halving process includes the modular exponentiation. Using the results, we perform a regression analysis to formulate mathematical equations for the cost in the section that follows.
@@ 5.1 Modular Exponentiation Cost
This part explores strategies for optimizing the computational cost of modular exponentiation, a frequent operation in various cryptographic algorithms. It discusses algorithms such as Square and Multiply [11], Dimitrov multiexponentiation [12], and the ModExp contract introduced by EIP-198.
5.1.1 Square and Multiply
The square and multiply algorithm, also called binary exponentiation, is an efficient approach for computing exponentiation operations $x^{n}$, particularly beneficial for large number modular exponentiation. As described in Algorithm 5, by converting the exponent $b$ into its binary representation, the algorithm optimizes computation through a series of square (squaring) and multiply (multiplying by the base $x$) operations based on the binary digits of $n$. For an exponent represented by $k$ bits, the algorithm requires at most $k-1$ squaring and $k-1$ multiplication operations, making the computational complexity $O(\log_{2}n)$. This optimization is simple but significantly enhances efficiency for large-scale modular exponentiation tasks.
Algorithm 5 Square and Multiply for Modular Exponentiation
1:input: $a,b,N$
2:output: $a^{b}$ mod $N$
3:$r\leftarrow 1$
4:$a\leftarrow a\mod N$
5:while $b>0$ do
6:     if $b\mod 2=1$ then
7:         $r\leftarrow(r\cdot a)\mod N$
8:     end if
9:     $b\leftarrow b>>1$
10:     $a\leftarrow(a\cdot a)\mod N$
11:end while
12:return $r$
5.1.2 Dimitrov Multiexponentiation
The Dimitrov algorithm [12, 3] presented in Algorithm 6 enhances the efficiency of large integer multiplication, crucial in fields such as cryptography and computational mathematics, by significantly reducing the computational complexity. Traditionally, the multiplication of large integers ($x^{a}y^{b}$) exhibits a complexity of $O(2\log_{2}max(a,b))$. Utilizing the binary exponentiation method [13], the Dimitrov algorithm achieves a computational complexity closer to $O(\frac{7}{4}\log_{2}max(a,b))$. This characteristic allows for the rapid and resource-efficient multiplication of very large integers, making the algorithm particularly valuable for applications requiring efficient large-scale numerical operations. In the generalized verification (Algorithm 4), lines 9 – 12 perform multiexponentiation only if $T/2$ is odd. Therefore, we considered the Dimitrov algorithm as a candidate algorithm to apply for optimization.
Algorithm 6 Dimitrov Multiexponentiation Algorithm
1:input: $x,y,a,b,N$
2:output: $x^{a}\cdot y^{b}\mod N$
3:$h\leftarrow\max(\lfloor\log_{2}a\rfloor+1,\lfloor\log_{2}b\rfloor+1)$
4:$z\leftarrow 1$
5:$q\leftarrow x\cdot y\mod N$
6:for $i=h-1$ down to 00 do
7:     $z\leftarrow z\cdot z\mod N$
8:     if $a_{i}=1$ and $b_{i}=0$ then
9:         $z\leftarrow z\cdot x\mod N$
10:     else if $a_{i}=0$ and $b_{i}=1$ then
11:         $z\leftarrow z\cdot y\mod N$
12:     else if $a_{i}=1$ and $b_{i}=1$ then
13:         $z\leftarrow z\cdot q\mod N$
14:     end if
15:end for
16:return $z$
Figure 3: Experiment results of the modular exponentiation cost
(a) Gas usage of exponentiation by squaring, square and multiply, and precompiled ModExp, x is 2048-bit level
(b) Gas usage of exponentiation by squaring, square and multiply, and precompiled ModExp, x is 3072-bit level
(c) Gas usage of Dimitrov multiexponentiation and only precompiled ModExp, x is 2048-bit level
(d) Gas usage of Dimitrov multiexponentiation and only precompiled ModExp, x is 3072-bit level
5.1.3 ModExp by EIP-198
EIP-198 [14], known for introducing "Big Integer Modular Exponentiation", introduces the ModExp (Modular Exponential) precompiled contract to the EVM, offering a streamlined and cost-efficient approach to conducting large integer modular exponentiation, a foundational operation in numerous cryptographic algorithms.
This function calculates $a^{b}\mod n$, where $a$, $b$, and $n$ are large numbers, and prior to EIP-198, such execution on-chain was not officially supported by any operations. The ModExp contract mitigates these expenses by providing a standardized computation method, thus facilitating the incorporation of cryptographic computations within smart contracts. Building on this, EIP-2565 [15] contributed to further refining the gas cost model for these modular exponentiation operations by proposing adjustments to the pricing formula.
5.1.4 Comparison of Exponentiation Techniques and Regression
Figure 3(a) and Figure 3(b) present the gas usage experiment results of three modular exponentiation techniques. We name a naive exponentiation approach as exponentiation by squaring in the figures. The comparing data named ‘Precompile’ is the gas usage when only the precompiled ModExp is used for multiexponentiation. Even if square and multiply (Algorithm 5) reduces the gas usage, it is obvious that the precompiled ModExp uses outstandingly less gas for the modular exponentiation than the others. Figure 3(c) and Figure 3(d) show the gas usage experiment results for modular multiexponentiation. We implemented Dimitrov multiexponentitaion including the precompiled ModExp. The latter takes significantly less gas than the Dimitrov algorithm. Therefore, we applied ModExp for both the modular exponentiation and multiexponentiation operations in the implementation in Algorithm 4.
In the experiment, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$. As Equation 1, we denote $\tau$ as the logarithmic value of T. We could get the following results from a simple regression process.
@@ 5.2 Halving Cost
The halving cost of the arithmetic operations in the verification indicates the gas usage for lines 5 – 13 in Algorithm 4. The modular exponentiation operations in lines 8, 10, and 12 use the precompiled ModExp since its best efficiency in the previous subsection. For simplicity, as we assumed $T=2^{\tau}$, in this case, the if statement in line 9 is always false, therefore line 10 is never executed. Hence, the gas usage will be clearly proportional to the number of the repetition time, that is, $\tau-\delta$.
Figure 4: Experiment results of the halving repetition cost
(a) Gas usage of the halving algorithm ($\lambda=2048$)
(b) Gas usage of the halving algorithm ($\lambda=3072$)
The results of our experiments demonstrate the anticipated correlation. Figure 4(a) and Figure 4(b) display the gas cost of the repeated halving operations with $\lambda=2048$ and $\lambda=3072$ correspondingly. The data exhibits a clear linear relationship, allowing us to perform a regression analysis and model the results using a linear function. By employing the method of least squares approximation, we obtain the subsequent functions as Equation 7 and 8:
$\mathcal{G}_{halving,2048}(\tau)=165482.26\cdot\tau-155642.3$ (7)
$\mathcal{G}_{halving,3072}(x)=306089.32\cdot\tau-238572.65$ (8)
@@ 6.1 Theoretical Analysis on Proof Generation for Gas Usage Optimization
The optimized proof generation in this research means minimizing the verification gas cost. We can define the total gas cost function $g_{total}$ using the previous regression results.
@@
Also, we denote the minimizing $\delta$ as
$\delta_{m}=\arg\min_{\delta}\mathcal{G}_{total}(\tau,\delta).$ (12)
From Theorem 1, we can get the unique $\delta_{m}$ optimizing the gas cost. Applying all the regression results in Equation 3, 4, 5, 6, 7, and 8, we get the optimizing $\delta_{m}\approx 9.47$ for $\lambda=2048$ and $\delta_{m}\approx 9.18$ for $\lambda=3072$ respectively. In practice, $\delta$ for implementation must be an integer. Therefore, we can get the practical integer $\delta$ with the following corollary.
Corollary 2.
  1. 1.
  2. 2.
  3. 3.
@@
6.2 Optimized Verification Gas Cost
Figures 5(a) and 5(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively. Consistent with the theoretical predictions in Theorem 1, the gas usage graph shows a decline to a minimal point before increasing sharply. Optimal gas costs were observed at $\delta=9$ for both configurations, aligning perfectly with the findings of Corollary 2. This alignment between theoretical expectations and empirical results confirms the accuracy of our model and highlights its potential for optimizing Pietrzak VDF verification in blockchain applications.
Table 1 details the gas consumption and calldata sizes for the implemented parameters, uniformly using $\delta=9$. Contrary to the previous expectation of calldata sizes around 40KB [3], the sizes recorded were significantly lower, ranging from 5.38 to 7.25 KB for $\lambda=2048$ and from 7.13 to 9.63 KB for $\lambda=3072$. Gas usage varied from 1.97 to 2.80 million for $\lambda=2048$ and from 3.68 to 5.25 million for $\lambda=3072$. These findings not only demonstrate the efficacy of our approach but also highlight potential areas for further optimization in VDF implementations. Such improvements could lead to significant cost reductions in blockchain networks, enhancing the overall efficiency of cryptographic operations.
Figure 5: Experiment results for halving proof skipping parameter $\delta$ with $\tau=20,21,22,23,24,25$
(a) Skipping parts of halving proofs ($\lambda=2048$)
(b) Skipping parts of halving proofs ($\lambda=3072$)
Table 1: Implementation results based on the analysis
$\lambda$ $\delta$ $\tau$ Gas used Calldata size (KB)
2048 9 20 1970844 5.38
21 2133372 5.75
22 2297629 6.13
23 2470594 6.50
24 2620907 6.89
25 2799805 7.25
3072 9 20 3680363 7.13
21 3967114 7.63
22 4318050 8.13
23 4614955 8.63
24 4947040 9.13
25 5254521 9.63
7 Discussion
This section discusses the implementation of VDFs on Ethereum, focusing on programming choices for Ethereum’s Virtual Machine and their impact on gas optimization. We introduce challenges associated with VDF deployment on Ethereum’s Layer 1 due to high costs and explore Layer 2 solutions. The discussion concludes with an analysis of economic factors critical for VDF integration and sustainability on the blockchain.
7.1 Implementation Languages
There are main development languages in EVM-compatible network environments, for example, Viper, Huff, Solidity, and Yul. For our implementation, which requires dynamically adjusting proof lengths, we excluded Vyper due to its inability to handle dynamic data structures efficiently. Instead, we chose Solidity combined with Yul, leveraging Solidity for its comprehensive features and developer familiarity, and Yul for its superior gas optimization capabilities. This combination provides a balance between high-level language safety and low-level control over the Ethereum Virtual Machine.
Considering future enhancements, we are exploring Huff, noted for its exceptional gas efficiency and direct control over the EVM. Huff originated from the Aztec team where contracts needed extreme efficiency, even more than what Yul could provide. Although Huff offers significant optimization opportunities, it lacks the built-in safety checks of higher-level languages, raising the implementation complexity. Nevertheless, given our implementation’s emphasis on gas optimization, the benefits of adopting Huff justify its consideration for future development.
7.2 Practicality of VDF on EVM
Implementing VDFs on Ethereum’s Layer 1 presents significant challenges primarily due to high gas costs and the computational intensity required. Our detailed optimization trials reveal that despite employing cutting-edge optimization techniques, verifying a Pietrzak VDF on Layer 1 still demands at least one million gas units. This high resource consumption makes VDF applications impractical for widespread use on Ethereum’s primary network, especially for operations requiring frequent and repetitive VDF verifications. The economic implications of these requirements can stifle innovation and discourage developers from incorporating VDFs into their decentralized applications, limiting the exploration of potentially valuable use cases such as randomized governance protocols or secure public randomness in blockchain networks.
However, the emergence of Layer 2 solutions, such as optimistic rollups and zkRollups, offers a promising alternative for leveraging VDFs within the Ethereum ecosystem. These technologies provide a secondary processing layer that can handle complex computations off the main Ethereum chain, thereby significantly reducing gas costs and improving transaction throughput. Platforms like Optimism capitalize on these advantages, offering gas prices that are often thousands of times cheaper than those on Layer 1. This dramatic cost reduction opens up new possibilities for using VDFs on Ethereum, allowing for more frequent and economically viable implementations. Layer 2 solutions not only mitigate the high cost barriers associated with Layer 1 but also broaden the scope for adopting VDFs in a variety of applications, thus encouraging further development and research into this promising area of blockchain technology.
7.3 Lack of Economic VDF Protocols in Ethereum
The integration of VDFs into smart contracts holds great promise for enhancing blockchain security and reliability. However, beyond the technical challenges, there exists a substantial hurdle in the lack of a robust economic model specifically designed to support the unique operational needs of VDFs, such as incentivizing both provers and verifiers and managing commit-reveal-recover protocols. The absence of such an economic framework complicates the practical deployment of VDFs as it does not adequately address compensation for participants who dedicate substantial computational resources to generate and verify proofs. This gap can deter the participation of potential contributors, ultimately hindering the scalability and effectiveness of VDFs in real-world blockchain applications.
Moreover, considering the decentralized essence of blockchain technology, it is crucial that any protocol, especially those as intricate as VDFs, is underpinned by an economic strategy that not only supports technical execution but also ensures economic sustainability without the need for central governance. This calls for innovative research into economic models that are specifically tailored to the decentralized context of blockchain networks. These models need to effectively balance incentives and deterrence, fostering a system where honest participation is rewarded while dishonest behavior is penalized. Such economic strategies are vital for maintaining the integrity and fairness of the blockchain ecosystem, ensuring that VDFs are not only technically feasible but also economically viable and beneficial to all participants, thereby aligning with the overarching principles of decentralization and trust minimization in blockchain systems.
7.4 Future Works on a General Environment
Through this research, we investigated the efficient implementation of Pietrzak VDF verification within the Ethereum Virtual Machine (EVM) environment and observed results that align both theoretically and experimentally. During our study, we identified significant effects of generalized proof generation and verification, which were also included within the predictable realms of theoretical analysis. We anticipate that such an approach could be applicable not only within the EVM but also in more general computing environments. Although resource pricing in general systems would differ from the gas costs in the EVM, with elements like the repetition of halving protocols and resource consumption proportional to the proof length being predictable, the costs of exponentiation in VDF ($x^{2^{2^{\tau}}}$) are expected to be exponential. Therefore, based on the outcomes of this study, further research on optimizing Pietrzak VDFs in general computing environments appears reasonable.
8 Conclusion
In conclusion, our research has successfully demonstrated the feasibility and significance of implementing VDFs within blockchain environments, particularly through smart contracts in Ethereum. By focusing on the Pietrzak VDF construction, we have not only bridged a crucial gap in the academic studies of VDFs but also highlighted their practical utility in enhancing the security and efficiency of blockchain applications. Contrary to previous expectations, our work represents the first academic result in the cost-effective on-chain verification of the Pietrzak VDF, showing significant potential in ensuring cryptographically secure random number generation across various blockchain applications. Through strategic implementation, we have achieved substantial cost reductions, making the Pietrzak VDF more accessible and applicable. This research opens new avenues for the integration of VDFs in blockchain systems, promising to enhance security, fairness, and functionality in the evolving distributed ledger technologies.

### v4

@@ 3.3 Environment for Gas Usage Measurement
3.3.1 Development Environment and Tools
The development of the contract was conducted using Solidity version 0.8.26 and EVM version Cancun. At the start of the development process, it was the latest compiler version fully supported by Foundry, which is a comprehensive Ethereum development toolchain that provides a robust environment for building, testing, and debugging smart contracts. Furthermore, gas optimization is best achieved with the latest version of the compiler.
@@ 3.3.2 Optimizer Configuration
The core parameter for the Solidity optimizer configuration is Runs which indicates how often each opcode will be used in the deployed contract. That is, higher Runs implies less gas cost for execution and higher gas cost for deployment. We set the optimizer parameter Runs to 4,294,967,295($2^{32}-1$), which is the maximum value of the parameter, indicating the highest level of optimization for execution.
@@
In this step, we could enhance the Yul optimizer’s efficiency by replacing the msize() operation with the free memory pointer. This adjustment necessitated manual updates to the memory indexes, a process rigorously tested to ensure the integrity of our implementation. We also applied memory-safe assembly blocks in our code. This change not only aligns with Solidity’s memory model but also activates Via-IR in the compiler settings, resulting in significant gas savings.
3.3.3 Gas Measurement in Local Testing Enviornment
As the test target code, we implemented contracts for each test in one main source file ensuring that each contract contains only one external function. This approach minimizes the error margin that can arise when multiple external functions are present, as the order of the function selectors impacts the gasUsed.
For the cost (gas usage) measurement, we utilized the Anvil node, provided by the Foundry toolchain, to create a local testing environment. The gas usage was obtained using the Forge testing framework provided by Foundry. The Forge standard library’s Test contract includes the lastCallGas() function, which retrieves the gasUsed in the last call. Additionally, isolate flag must be set when running the Forge’s test command to ensure all top-level calls are executed as separate transactions in distinct EVM contexts. This setup allows for more precise gas accounting. It is important to note that the actual gas cost incurred on the real network is determined by multiplying the estimated gasUsed by the gasPrice at the Ethereum network circumstance.
4 Cost measurement and approximation
@@ 4.1 Data Structure
In this study, the VDFs are implemented with key sizes of 2048 and 3072 bits. However, the EVM natively supports arithmetic operators only up to 256-bit numbers. To facilitate efficient computation of numbers exceeding 256 bits, it is necessary to employ customized data formats. We revised the big number library implementation for Solidity developed by Firo [12]. The original data structure in this library includes bytes data (bytes val), bit length (uint bitlen), and a sign indicator (bool neg). Among the three elements, we deleted the sign indicator and its related low-level logic since our implementation does not require handling negative values. This library optimizes on-chain operations such as addition and subtraction by adding leading zeros for 32-byte alignment with the EVM’s memory word size, eliminating offset management. Also, it utilizes the bit length data field to optimize multiplication and comparison operations. Therefore, we format the bytes data and pre-calculate bit length off-chain before submission, significantly reducing unnecessary on-chain computations. For example, the data format for a 2048-bit representation in JSON is as shown in Listing 1. In this way, we can reduce the gas cost even though the calldata size increases. This decision is reasonable because the primary goal is to reduce the gas cost, not the transaction calldata size.
@@
4.2 Gas Cost for Calldata and Dispatch
The gas cost for Ethereum transaction calldata, as described in EIP-2028 "Transaction data gas cost reduction" [13], is divided into two cases: 16 gas per non-zero byte and 4 gas per zero byte.
@@ 4.2 Gas Cost for Calldata and Dispatch
In our implementation, the function verifyRecursiveHalvingProof processes the refined verification, Algorithm 2. A call to this function involves encoding parameter values. The process starts by computing the keccak256 of the UTF-8 byte representation of the string111verifyRecursiveHalvingProof((bytes,uint256)[],(bytes,uint256),(bytes,uint256),(bytes,uint256),uint256,uint256), then using the first four bytes, which results in 0xd8e6ac60.
@@
Considering the prediction of data and its cost, we performed experiments to test the size of calldata and the related gas usage. The related gas usage comprises of the intrinsic gas and data dispatching gas. The intrinsic gas in EVM refers the minimum cost for a transaction. It contains the constant transaction cost (21000 gas) and the cost for data supply. For verification, we need not only to allocate data in a transaction, but also to dispatch data into the verification function. The dispatching cost is proportional to the data size. Therefore, we decided to include the dispatching cost into the data cost in this study. Figure 1(a) for $\lambda=2048$ and 1(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage. We can see the gas usage and the calldata size simply and linearly increase with the number of proofs.
Figure 1: Experiment results of the calldata and its dispatch cost
(a) $\lambda=2048$
(b) $\lambda=3072$
Figure 2: Experiment results of the halving algorithm cost
(a) $\lambda=2048$
(b) $\lambda=3072$
Figure 3: Experiment results of the ModExp cost
(a) $\lambda=2048$
(b) $\lambda=3072$
Finally, to describe the relationship between transaction size and gas costs, we performed a basic regression analysis. By applying the least squares method to approximate the best fit, we derive the following linear regression equations 1 and 2 for data sizes of 2048 and 3072 bytes, respectively:
@@ 4.3 Modular Exponentiation Cost
EIP-198 [14], known for "Big Integer Modular Exponentiation", introduced the ModExp (Modular Exponential) precompiled contract to the EVM, offering a streamlined and cost-efficient approach for large integer modular exponentiation.  [^2] EIP-2565 [15] contributed to further refining the gas cost model for these modular exponentiation operations by proposing adjustments to the pricing formula. As its obvious that a precompile provides the best efficinecy, we applied ModExp for both the modular exponentiation and multi-exponentiation operations in the implementation in Algorithm 2. In Figure 3(a) and Figure 3(b) illustrating the experiment results, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$, where we denote $\tau$ as the logarithmic value of $T$. We could get the following results from a simple regression process.
@@ 4.4 Halving Cost
The halving cost of the arithmetic operations in the verification indicates the gas usage for lines 5 – 13 in Algorithm 2. The modular exponentiation operations in lines 8, 10, and 12 use the precompiled ModExp. For simplicity, as we assumed $T=2^{\tau}$, in this case, the if statement in line 9 is always false, therefore line 10 is never executed. Hence, the gas usage is clearly proportional to the number of the repetition time, that is, $\tau-\delta$.
The results of our experiments demonstrate the anticipated correlation. Figure 2(a) and Figure 2(b) display the gas cost of the repeated halving operations with $\lambda=2048$ and $\lambda=3072$ correspondingly. The data exhibits a clear linear relationship, allowing us to perform a regression analysis and model the results using a linear function. By employing the method of least squares approximation, we obtain the subsequent functions as Equation 5 and 6:
$\mathcal{G}_{halving,2048}(\tau)=160604.41\cdot\tau-157178.66$ (5)
$\mathcal{G}_{halving,3072}(x)=299896.98\cdot\tau-248119.67$ (6)
@@ 5.1 Theoretical Analysis on Proof Generation for Gas Usage Optimization
The optimized proof generation in this research means minimizing the verification gas cost in EVM. We can define the total gas cost function $g_{total}$ using the previous regression results.
@@
Proof.
See Appendix B. ∎
Also, we denote the minimizing $\delta$ as
$\delta_{m}=\arg\min_{\delta}\mathcal{G}_{total}(\tau,\delta).$ (10)
From Theorem 1, we can get the unique $\delta_{m}$ optimizing the gas cost. Applying all the regression results in Equation 1, 2, 3, 4, 5, and 6, we get the optimizing $\delta_{m}\approx 9.43$ for $\lambda=2048$ and $\delta_{m}\approx 9.15$ for $\lambda=3072$ respectively. In practice, $\delta$ for implementation must be an integer. Therefore, we can get the practical integer $\delta$ with the following corollary.
Corollary 2.
  1. 1.
  2. 2.
  3. 3.
Proof.
See Appendix C. ∎
5.2 Optimized Verification Gas Cost
Figures 4(a) and 4(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively. Consistent with the theoretical predictions in Theorem 1, the gas usage graph shows a decline to a minimal point before increasing sharply. Optimal gas costs were observed at $\delta=9$ for both configurations, aligning perfectly with the findings of Corollary 2. This alignment between theoretical expectations and empirical results confirms the accuracy of our model and highlights its potential for optimizing Pietrzak VDF verification in blockchain applications.
[…]
Figure 4: Results for halving proof skipping parameter $\delta$ with $\tau=20,21,22,23,24,25$ for $y=x^{2^{2^{\tau}}}$
(a) Skipping parts of halving proofs ($\lambda=2048$)
(b) Skipping parts of halving proofs ($\lambda=3072$)
Table 1: Implementation results based on the analysis
$\lambda$ $\delta$ $\tau$ Gas used Calldata size (KB)
2048 9 20 1817855 5.38
21 1973222 5.75
22 2129116 6.13
23 2286058 6.50
24 2445803 6.88
25 2603753 7.25
3072 9 20 3495396 7.13
21 3776893 7.63
22 4078468 8.13
23 4362630 8.63
24 4658937 9.13
25 4958138 9.63
6 Conlusion and Future Works
Our study demonstrates that Pietrzak’s VDF can be effectively optimized for the EVM, reducing gas costs from 4M to 2M and minimizing proof size to under 8 KB with a 2048-bit RSA key. These optimizations validate the theoretical models and show practical viability. While our paper does not address the Wesolowski VDF implementation, which poses challenges due to its hash-to-prime function, exploring indirect or approximate implementation methods could make it feasible for the EVM. Future work should focus on utilizing Wesolowski VDF verification on EVM as it has a big benefit on the constant and short proof length and verification cost. We plan to implement on-chain Wesolowski VDF verification both directly and indirectly. For indirect implementation, a hash-to-prime function could be executed in a Truebit-style proof-of-fraud system. For direct implementation, refining the hash-to-prime function to be viable and secure on the EVM is essential.
Appendix A Code availability
To facilitate reproducibility and peer review, we have made the source code for our implementation available anonymously. The repository contains all the scripts and supplementary materials necessary to reproduce the results discussed in this paper. You can access the repository at the following link: https://github.com/usgeeus/Pietrzak-VDF-solidity-verifier
Appendix B Proof of Theorem 1
@@
Appendix C Proof of Corollary 2